# CoVLA: Comprehensive Vision-Language-Action Dataset for Autonomous Driving

### v1

@@ 2.1 Datasets for Autonomous Driving
2.2 Multi-modal LLMs for Driving Tasks
Addressing the complex and often unpredictable nature of real-world driving scenarios, particularly those falling within the “long tail” of the driving distribution (e.g., navigating road closures), necessitates a higher level of reasoning and decision-making capabilities in autonomous systems. Recent research has increasingly focused on incorporating Multi-modal Large Language Models (MLLMs) into autonomous driving systems, aiming to leverage their powerful reasoning and language understanding capabilities to address these challenges. For instance, ADAPT [24] utilizes Vision-Language Models (VLMs) to generate natural language explanations of driving actions and provide reasoning behind them, enhancing the interpretability of the system. Early attempts to predict driving actions using language models involved employing GPT-3.5 and GPT-4 [1] to predict high-level driving commands from situational descriptions [20, 54]. These initial efforts set the stage for more recent studies, which have explored fine-tuning LLMs to predict various types of driving behavior like trajectories [57, 49] or control signals such as steering and acceleration [10, 37, 47] through textual outputs. While these approaches handle driving maneuver directly through language, LMDrive [48] and DriveMLM [53] predict control signals by passing the embeddings from the Transformer outputs to decoders. CarLLaVA [45] uses the LLaVA architecture to predict trajectories by taking images, vehicle speed, and target location coordinates as inputs. Notably, this approach has achieved significantly superior performance compared to existing methods on the CARLA leaderboard 2.0, demonstrating the effectiveness of VLMs for complex trajectory prediction tasks.
@@ 3.1.1 Raw Data
We develop multiple data-collection vehicles to acquire real-world driving data, including sensor data from the front-facing camera, Controller Area Network (CAN) bus, Global Navigation Satellite System (GNSS), and Inertial Measurement Unit (IMU) and deploy them around Tokyo, Japan.
  • Data collection period: 6 months (summer to winter), resulting in over 1000 hours of raw data
  • Data collection environments: Various locations in and around Tokyo (urban centers, complex highway interchanges, narrow residential streets, mountainous areas with winding roads), different weather conditions (sunny, cloudy, rainy, heavy rain), and various times of day (daytime, evening, nighttime)
  • Equipment: Front-facing camera (1928$\times$1208 pixels, 20 FPS, H.265 codec), CAN bus (accelerator/brake pedal position, steering wheel angle, turn signal status, gear position, vehicle speed), GNSS and IMU (binary format with timestamps), in-vehicle storage devices
@@ 3.1.3 Auto Labeling
Figure 3: Sample frames from CoVLA-Dataset. Estimated trajectories (green line) and captions generated by the captioner model are shown. The key objects are highlighted in blue bold text, while the failures in captions are shown in red bold text.
@@
To acquire future trajectory data, which will be predicted by the VLA models, we estimate the vehicle’s travel path with a Kalman Filter [25] using GNSS and IMU sensors. For each timestamp, we annotate future trajectory data points for the subsequent three seconds (60 frames). The trajectory data is expressed in the global coordinate system centered on the data collection vehicle. In some instances, GNSS data instability occasionally leads to erroneous trajectories characterized by significant vibrations. We implement a heuristic approach to identify and remove these inaccurate trajectories from the dataset.
@@ 3.1.4 Auto Captioning
To mitigate this issue, we use comprehensive rule-based captions as factual constraints. These captions are then provided as context to the VLM, along with a prompt instructing it to supplement the captions with any additional information not already covered by the rules. This supplementary information, derived by querying the model and examining its internal token probability distributions, includes details regarding road type (e.g., narrow/wide, highway/non-highway, tunnel/non-tunnel), weather conditions (e.g., sunny, cloudy, rainy), potential risks (e.g., pedestrian presence), and other relevant factors. For instance, to determine the weather, we calculate the probability of each weather-related token (e.g., “sunny”, “cloudy”, or “rainy”) given the query “What is the weather in this video?” and select the token with the highest probability. Using these augmented rule-based captions as factual anchors, we prompt the VLM to generate free-form captions, focusing on identifying potential risks within the driving environment.
@@ 3.2 Analysis
To assess the impact of our scene sampling methodology described in Section §3.1.2, we first analyze the distribution of key driving parameters. As shown in 4(a) and 4(b), the distributions of vehicle speeds and steering angles become more uniform after sampling, with a noticeable increase in lower speed data points and a diminished central peak around zero degrees for steering angles, leading to a more balanced representation across different ranges. Consequently, CoVLA-Dataset exhibits diverse driving maneuvers, with active turn signals in 16.11% of frames and traffic lights in 22.90%. This diversity, including starting/stopping, turning, and lane changing, emphasizes the need for robust perception, decision-making, and control algorithms in autonomous driving systems.
@@
CoVLA-Dataset distinguishes itself from other datasets through its seamless integration of vision, language, and action modalities. This integration is a result of our focus on synchronized real-world data acquisition from various sensors, coupled with our automated labeling and captioning pipeline. This comprehensive annotation approach captures all relevant aspects of the driving environment, making it a valuable resource for training and evaluating autonomous driving systems.
@@
Figure 3 presents sample frames from CoVLA-Dataset, highlighting both successful examples and instances where auto-captioning encountered challenges. While our proposed method effectively enables scalable and descriptive video captioning, analysis of the generated captions reveals areas for further refinement. We identify two primary categories of errors. Firstly, the captioning model occasionally exhibits object hallucination, describing objects absent from the scene, such as a “wooden fence.” Furthermore, the VLM occasionally misidentifies object locations and scene elements, such as incorrectly labeling left-side objects as “right side.” Another limitation we reveal is the recognition of local landmarks, particularly those unique to Japan. For instance, Japanese traffic signs often have unique designs and meanings, such as including Japanese characters, that necessitate proper identification and interpretation.
@@ 4.1 Model
CoVLA-Agent as illustrated in Figure 5, is a VLA model designed for autonomous driving. We utilize the pretrained Llama-2 (7B) [52] as a language model and CLIP ViT-L (224×224 pixels) [43] as a vision encoder. In addition, our model takes the ego vehicle’s speed as an input, which is transformed into an embedding vector using a Multi Layer Perceptron (MLP). The visual features extracted by CLIP ViT-L are concatenated with the speed embedding and text embedding, and then fed into the Llama-2 model. For trajectory prediction, special tokens are used as trajectory queries. The outputs of these trajectory queries are processed by an MLP layer, resulting in a sequence of 10 $(x,y,z)$ coordinates representing the predicted trajectory of the vehicle relative to its current position, spanning a three-second time horizon.
@@ 4.4 Discussion
6(d) shows the results of trajectory prediction at an intersection. In this scene, the ground truth caption includes the phrase “moving straight” and the trajectory predicted with that condition is straight. In contrast, the predicted trajectory based on the predicted caption including phrase “turning right”, accurately reflects a right-turn trajectory. While this result indicates the difficulty of estimating intention from a single frame in such scenes, it also underscores the consistency between language and action modality in CoVLA-Agent.
To examine the impact of these captioning errors on trajectory, we investigate how errors in predicted captions affect ADE and FDE. Specifically, we analyze the mean ADE and FDE values for each word in the captions when the word appears in the ground truth rule-based caption part but not in the predicted caption, and vice versa. Excluding stop words, the top ten words with frequencies exceeding ten that led to the largest ADE and FDE values are shown in Table 3, though their rankings varies between the two metrics. Large ADE and FDE occur when there are discrepancies in the predicted caption, particularly with words related to vehicle motion direction and acceleration, such as “deceleration”, “left”, “curve”, and “turning”. Based on these observations, we attribute the relatively low performance in the trajectory predicted using the generated caption largely to the difficulty of estimating intention from a single frame.

### v2

@@ 2.1 Datasets for Autonomous Driving
2.2 MLLMs for Driving Tasks
Addressing the complex and often unpredictable nature of real-world driving scenarios, particularly those falling within the “long tail” of the driving distribution (e.g., navigating road closures), necessitates a higher level of reasoning and decision-making capabilities in autonomous systems. Recent research has increasingly focused on incorporating MLLMs into autonomous driving systems, aiming to leverage their powerful reasoning and language understanding capabilities to address these challenges. For instance, ADAPT [24] utilizes Vision-Language Models (VLMs) to generate natural language explanations of driving actions and provide reasoning behind them, enhancing the interpretability of the system. Early attempts to predict driving actions using language models involved employing GPT-3.5 and GPT-4 [1] to predict high-level driving commands from situational descriptions [20, 54]. These initial efforts set the stage for more recent studies, which have explored fine-tuning LLMs to predict various types of driving behavior like trajectories [57, 49] or control signals such as steering and acceleration [10, 37, 47] through textual outputs. While these approaches handle driving maneuver directly through language, LMDrive [48] and DriveMLM [53] predict control signals by passing the embeddings from the Transformer outputs to decoders. CarLLaVA [45] uses the LLaVA architecture to predict trajectories by taking images, vehicle speed, and target location coordinates as inputs. Notably, this approach has achieved significantly superior performance compared to existing methods on the CARLA leaderboard 2.0, demonstrating the effectiveness of VLMs for complex trajectory prediction tasks.
@@ 3.1.1 Raw Data
We develop multiple data collection vehicles to acquire real-world driving data, including sensor data from the front-facing camera, Controller Area Network (CAN) bus, Global Navigation Satellite System (GNSS), and Inertial Measurement Unit (IMU) and deploy them around Tokyo, Japan.
  • Data collection period: 6 months (summer to winter), resulting in over 1000 hours of raw data.
  • Data collection environments: Various locations in and around Tokyo (urban centers, complex highway interchanges, narrow residential streets, mountainous areas with winding roads), different weather conditions (sunny, cloudy, rainy, heavy rain), and various times of day (daytime, evening, nighttime).
  • Equipment: Front-facing camera (1928$\times$1208 pixels, 20 FPS, H.265 codec), CAN bus (accelerator/brake pedal position, steering wheel angle, turn signal status, gear position, vehicle speed), GNSS and IMU (binary format with timestamps), in-vehicle storage devices.
@@ 3.1.3 Auto Labeling
Figure 3: Frame examples from CoVLA-Dataset. Estimated trajectories (green line) and captions generated by the captioner model are shown. The key objects are highlighted in blue bold text, while the failures in captions are shown in red bold text.
@@
To acquire future trajectory data, which will be predicted by the VLA models, we estimate the vehicle’s travel path with a Kalman Filter [25] using GNSS and IMU sensors. For each timestamp, we annotate future trajectory data points for the subsequent three seconds (60 frames). The trajectory data is expressed in the global coordinate system centered on the data collection vehicle. In some instances, GNSS data instability occasionally leads to erroneous trajectories characterized by significant vibrations. We implement a heuristic approach to identify and remove these inaccurate trajectories from the dataset. Further details are reported in the supplementary material.
@@ 3.1.4 Auto Captioning
To mitigate this issue, we use comprehensive rule-based captions as factual constraints. These captions are then provided as context to the VLM, along with a prompt instructing it to supplement the captions with any additional information not already covered by the rules. This supplementary information, derived by querying the model and examining its internal token probability distributions, includes details regarding road type (e.g., narrow/wide, highway/non-highway, tunnel/non-tunnel), weather conditions (e.g., sunny, cloudy, rainy), potential risks (e.g., pedestrian presence), and other relevant factors. For instance, to determine the weather, we calculate the probability of each weather-related token (e.g., “sunny,” “cloudy,” or “rainy”) given the query “What is the weather in this video?” and select the token with the highest probability. Using these augmented rule-based captions as factual anchors, we prompt the VLM to generate free-form captions, focusing on identifying potential risks within the driving environment.
@@ 3.2 Analysis
To assess the impact of our scene sampling methodology described in Section 3.1.2, we first analyze the distribution of key driving parameters. As shown in 4(a) and 4(b), the distributions of vehicle speeds and steering angles become more uniform after sampling, with a noticeable increase in lower speed data points and a diminished central peak around zero degrees for steering angles, leading to a more balanced representation across different ranges. Consequently, CoVLA-Dataset exhibits diverse driving maneuvers, with active turn signals in 16.11% of frames and traffic lights in 22.90%. This diversity, including starting/stopping, turning, and lane changing, emphasizes the need for robust perception, decision-making, and control algorithms in autonomous driving systems.
@@
CoVLA-Dataset distinguishes itself from other datasets through its seamless integration of vision, language, and action modalities. This integration results from synchronized real-world data acquisition from various sensors, coupled with our automated labeling and captioning pipeline. This comprehensive annotation captures all relevant aspects of the driving environment, making it a valuable resource for training and evaluating autonomous driving systems.
@@
Figure 3 presents frames sampled from CoVLA-Dataset, highlighting both successful examples and instances where auto-captioning encountered challenges. While our proposed method effectively enables scalable and detailed video captioning, analysis of the generated captions reveals areas for further refinement.
We identify two primary categories of errors. Firstly, the captioning model occasionally exhibits object hallucination, describing objects absent from the scene, such as a “wooden fence.” Furthermore, the VLM occasionally misidentifies object locations and scene elements, such as incorrectly labeling left-side objects as “right-side.”
Another limitation we reveal is the recognition of local landmarks, particularly those unique to Japan. For instance, Japanese traffic signs often have unique designs and meanings, such as including Japanese characters, that necessitate proper identification and interpretation.
@@ 4.1 Model
CoVLA-Agent as illustrated in Figure 5, is a VLA model designed for autonomous driving. We utilize the pretrained Llama-2 (7B) [52] as a language model and CLIP ViT-L (224$\times$224 pixels) [43] as a vision encoder. In addition, our model takes the ego vehicle’s speed as an input, which is transformed into an embedding vector using a Multi Layer Perceptron (MLP). The visual features extracted by CLIP ViT-L are concatenated with the speed embedding and text embedding, and then fed into the Llama-2 model. For trajectory prediction, special tokens are used as trajectory queries. The outputs of these trajectory queries are processed by an MLP layer, resulting in a sequence of 10 $(x,y,z)$ coordinates representing the predicted trajectory of the vehicle relative to its current position, spanning a three-second time horizon.
@@ 4.4 Discussion
6(d) shows the results of trajectory prediction at an intersection. In this scene, the ground truth caption includes the phrase “moving straight” and the trajectory predicted with that condition is straight. In contrast, the predicted trajectory based on the predicted caption including phrase “turning right,” accurately reflects a right-turn trajectory. While this result indicates the difficulty of estimating intention from a single frame in such scenes, it also underscores the consistency between language and action modality in CoVLA-Agent.
To examine the impact of these captioning errors on trajectory, we investigate how errors in predicted captions affect ADE and FDE. Specifically, we analyze the mean ADE and FDE values for each word in the captions when the word appears in the ground truth rule-based caption part but not in the predicted caption, and vice versa. Excluding stop words, the top ten words with frequencies exceeding ten that led to the largest ADE and FDE values are shown in Table 3, though their rankings vary between the two metrics. Large ADE and FDE occur when there are discrepancies in the predicted caption, particularly with words related to vehicle motion direction and acceleration, such as “deceleration,” “left,” “curve,” and “turning.” Based on these observations, we attribute the relatively low performance in the trajectory predicted using the generated caption largely to the difficulty of estimating intention from a single frame.
@@ Acknowledgments
Appendix A Heuristic Trajectory Filtering
In some instances, trajectory data instability was detected. Specifically, we identified two erroneous behaviors:
  1. 1.
    Significant jumps
  2. 2.
    Movement in the wrong direction
To detect significant jumps, we filtered the trajectory data by the distance between adjacent points. Given a recording frequency of 20 Hz and a maximum speed of 100 km/h, the distance between points should be at most 1.38 meters, which is calculated as following:
$\frac{100\,\text{km/h}\times 1000\,\text{m/km}}{3600\,\text{s/h}\times 20}=1.38\,\text{m}$ (1)
With a tolerance rate of 1.15, the threshold was set to 1.59 meters. All trajectories exceeding this threshold were filtered.
To detect movement in the wrong direction, we manually checked 400 samples from all scenes and identified 43 invalid trajectories (10.75%). Observations revealed a vibration frequency of 10 Hz in these trajectories. To implement vibration detection, we smoothed the trajectory using a 3-point moving average and calculated the difference between the smoothed trajectory and the original trajectory. We then analyzed the variance of these differences. If the variance exceeded a certain threshold, the trajectory was classified as invalid. This method yielded a precision of 0.64 and a recall of 0.75 on the test dataset, reducing the invalid trajectory rate to 2.6%. Although this method has a relatively high false-positive rate, it is acceptable for the dataset’s scale.
Appendix B Privacy Protection for Dataset Publication
We published CoVLA-Dataset  [^3] https://huggingface.co/datasets/turing-motors/CoVLA-Dataset on HuggingFace. For privacy protection, we anonymized human faces and license plates in the CoVLA-Dataset images and videos, using Dashcam Anonymizer  [^4] https://github.com/varungupta31/dashcam_anonymizer.
Appendix C Example data in CoVLA-Dataset
We present sample data from CoVLA-Dataset. Each data point includes an image, a caption, and vehicle states. An example of an actual caption is shown in Figure 7. For auto-captioning, we use the traffic light detection provided by OpenLenda-s and the front car detection results from the sensor fusion. The list of vehicle states is indicated in Table 4. Each frame contains information on the speed, steering angle, coordinates of the trajectory, and more.
Figure 7: Example of image and caption in CoVLA-Dataset.
Key Value
frame_id 569
image_path images/2022-07-08–11-37-27–5_first/0569.png
vEgo 7.43082332611084
vEgoRaw 7.4395833015441895
aEgo 0.6044138669967651
steeringAngleDeg 0.6073870658874512
steeringTorque 69.0
brake 0.0
brakePressed false
gas 0.20499999821186066
gasPressed true
doorOpen false
seatbeltUnlatched false
gearShifter drive
leftBlinker false
rightBlinker false
orientations_calib $\begin{bmatrix}2.9467956252753775&0.9174552319868815&2.2181786819453384\end{bmatrix}$
orientations_ecef $\begin{bmatrix}2.9243210649189977&0.9224135550861058&2.1900513923432348\end{bmatrix}$
orientations_ned $\begin{bmatrix}-0.013463193567253392&0.006326533926443111&-2.990125370637735\end{bmatrix}$
positions_ecef $\begin{bmatrix}-3959574.486029379&3328427.354910454&3719065.7393601397\end{bmatrix}$
velocities_calib $\begin{bmatrix}7.317097759615114&0.003242519329502727&0.005369323447773883\end{bmatrix}$
velocities_ecef $\begin{bmatrix}-2.6767882666706004&3.547338396353873&-5.813015899212604\end{bmatrix}$
accelerations_calib $\begin{bmatrix}0.4734579094803297&0.08559864698994124&-0.13132037594775653\end{bmatrix}$
accelerations_device $\begin{bmatrix}0.4736293760658642&0.07819260264673351&-0.13526157094253702\end{bmatrix}$
angular_velocities_calib $\begin{bmatrix}0.011550541795216845&0.012243857869171634&-0.007753300486330907\end{bmatrix}$
angular_velocities_device $\begin{bmatrix}0.011675888068301523&0.01206267096884485&-0.007848971055415363\end{bmatrix}$
timestamp 1657248173200
extrinsic_matrix $\begin{bmatrix}-0.015688330416257182&-0.9998769191404183&0.00012959444326649344&0.0\\ -0.008260370686184616&2.879912020664621e-21&-0.9999658837914467&1.2200000286102295\\ 0.9998428078989188&-0.01568886620613436&-0.008259354077745229&0.0\\ 0.0&0.0&0.0&1.0\end{bmatrix}$
intrinsic_matrix $\begin{bmatrix}2648.0&0.0&964.0\\ 0.0&2648.0&604.0\\ 0.0&0.0&1.0\end{bmatrix}$
trajectory_count 60
trajectory $\begin{bmatrix}0.0&-0.0&0.0\\ ...&...&...\\ 2.2367286927600376&0.021314714278469867&0.011369742248091336\\ ...&...&...\\ 4.529136516056675&0.05988551136966269&0.03572205827208557\\ ...&...&...\\ 6.919731941586808&0.08958899605082428&0.050459818682157966\\ ...&...&...\\ 9.301852576186251&0.1417775525229876&0.07520389298492622\\ ...&...&...\\ 11.677091075613516&0.2108797042962825&0.09371664992653018\\ ...&...&...\\ 14.027571172292527&0.28508464282229706&0.10760938523286476\\ ...&...&...\\ 16.350994760951718&0.38061347529509826&0.1208462683905208\\ ...&...&...\\ 18.675065252972097&0.4777731491748536&0.14622174266379429\\ ...&...&...\\ 20.998125620182435&0.5860796022647292&0.1657656398148593\end{bmatrix}$
Table 4: Example of Vehicle States List.